SHOW ALL WORK. PROBLEMS 1 and 2 ARE CONNECTED

1. A major car manufacturing firm issues a 20 year $1,000,000 bond at par. The bond pays a 6% (APR) semiannual coupon. 5 years later, the Federal Reserve Board cuts the fed funds rate, causing interest rates for similar firms to fall to 4% (APR). What is the new bond price?

A. $1,222,368

B. $1,223,965

C. $1,000,000

D. $1,273,555

2.  . If you bought the bond at issue and held it to maturity, what is the effective annual rate (EAR) that you earned?

A. 6%

B. 6.09%

C. 4%

D. 4.04%
Answer #1

1)Here they are asking to find the amount received when they sold bonds , the formulae is below
Price=(coupon*(1-((1+i)^-n))/i)+(issue price*(1+i)^-n)
Coupon=maturity value*coupon rate
=1000000*(6%/2)=30000
n=15*2(since 5 years passed)
i=4%/2
issue price=1000000
substituting in formuale we get it as 1223965
it is option B

2)since here we are purchasing bond at issue price which is par and holding till maturity so our return is same as coupon rate and it is 6%
it is option A